OpenStudy (anonymous):

Help please! Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8>

OpenStudy (rational):

To check whether they are parallel or not, take the ratio of corresponding components : \[\dfrac{1}{-4} \\~\\~\\~\\ \dfrac{-2}{8}\] area they equal ?

OpenStudy (anonymous):

no

OpenStudy (rational):

Are you really sure ?

OpenStudy (rational):

\[\large \dfrac{1}{-4}~~~\stackrel{?}{=}~~~\dfrac{-2}{8}\]

OpenStudy (anonymous):

Its true sorry.

OpenStudy (rational):

Yes so the given vectors are parallel

OpenStudy (anonymous):

I didn't think it would be that simple...Thank you!

OpenStudy (rational):

yw

OpenStudy (anonymous):

a picture is worth a thousand words draw it it's a straight line